OpenStudy (anonymous):

The area of a rectangle is 12 square inches. The length is 5 more than twice the width. Find the length and the width.

OpenStudy (anonymous):

You wanna try write out the equation?

OpenStudy (anonymous):

F(12)=5+x ?

OpenStudy (anonymous):

Not quite :D

OpenStudy (anonymous):

So you have from the text that: length = 2*width+5 length*width=12

OpenStudy (anonymous):

12(2)+5?

OpenStudy (anonymous):

Do you know how to solve quadratic equations?

OpenStudy (anonymous):

3.5x2=7+5=12 ?

OpenStudy (anonymous):

Do you?? This question was a bit longer than the previous one

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I solved it using quadratic equation, but if don't you know that this exercise might be a bit hard??

OpenStudy (anonymous):

length = l, width = w l = 2w+5 (1.) l*w = 12 (2.) (2.) => l = 12/w plugging into (1.) gives 12/w = 2w+5 12 = w(2w+5) 12 = 2w^2+5w 6 = w^2+2.5w w=1.5 or -4 (but -4 is not possible since it's a length) l = 2*1.5+5 = 8 So length = 8, width = 1.5

OpenStudy (anonymous):

Other than that, I think you just have to guess something that works...

OpenStudy (anonymous):

ok i think your steps will help it is starting to come back i have not done this type work for about 10 years

OpenStudy (anonymous):

10 years is a loooooooong time :D

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

To clarify, I expressed length in width (l=12/w), so that I could solve for width. After simplifying, I got w^2+2.5w-6, which is a quadratic equation I solved using the quadratic formula. Then I could solve for l.

OpenStudy (anonymous):

ok thank you i have one more for you.

OpenStudy (anonymous):

The area of a rectangle is 28 square inches. The length is 8 more than thrice the width. Find the length and the width.

OpenStudy (anonymous):

That's pretty much the same exercise, just with other numbers

OpenStudy (anonymous):

28w=3w=8

OpenStudy (anonymous):

28w=3w+8 sorry

OpenStudy (anonymous):

It's not that simple, I'm afraid. You have to use the same method I did solving the other one.

OpenStudy (anonymous):

This is how you interpret the text, though:

OpenStudy (anonymous):

"The area of a rectangle is 28 square inches." - Area equals length multiplied by width. So A = l*w. This means 28 = l*w. "The length is 8 more than thrice the width. Find the length and the width." - Length is 8 more than 3*width. So l = 8+3*w. You get the two equations: l*w=28 l=8+3w

OpenStudy (anonymous):

Huh, all I was typing just disappeared :(

OpenStudy (anonymous):

oh that sucks i am sorry for being an inconveniance

OpenStudy (anonymous):

You can solve it with the same method I did though if you just follow step-by-step and just keep track of the differing numbers

OpenStudy (anonymous):

But I have to leave now, good luck though

OpenStudy (anonymous):

thanks